# The Influence of Muons, Pions, and Trapped Neutrinos on Neutron Star Mergers

### v1

@@ III.4 Gravitational Wave Emission
One of the main promising routes of obtaining additional information from the post-merger gravitational wave signal is via the dominant frequency peak of the gravitational wave strain Chatziioannou et al. (2017); Wijngaarden et al. (2022); Criswell et al. (2023); Breschi et al. (2022, 2023). In Fig. 14 we show the frequency spectrum of the post-merger gravitational wave strain, focusing mainly on the dominant frequency peak. We begin by comparing the impact of adding different microphysical correction to the equation of state. In line with previous works, we find that the adding of trapped electron neutrinos does not appreciable alter the gravitational wave emission Zappa et al. (2022); Most et al. (2024); Hammond et al. (2021), although current simulations are likely not in a convergent regime Espino et al. (2024); Most et al. (2024). We investigate the impact of adding $\pi$ to the EOS. We already found earlier that pions lead to a net softening (more compact neutron stars) of the equation of state, which generically should translate to higher frequencies. In line with this expectation, we find that for both EOS adding pions causes a shift of around $100\,\rm Hz$ of the post-merger gravitational wave signal, although the impact on SFHo seems slightly more pronounced. Such difference is consistent with uncertainties of quasi-universal relations Vretinaris et al. (2020); Topolski et al. (2024). Finally, we also investigate the impact of $\mu$, and trapped $\nu_{\mu}$. Similar to the addition of trapped $\nu_{e}$, we find that the impact on the post-merger gravitational wave signal is small, though we caution that $\mu$ can in principle regulate the population of $\pi$, as seen in the amount of pion condensation (Fig. 5). While these frequencies are out of the current frequency sensitivity bands for the LIGO-Virgo-KAGRA detectors Barsotti et al. (2010), these calculations stress the importance of capturing proper microphysics for accurate predictions of future NSNS waveforms to optimize future high frequency GW detectors.
@@ IV Conclusions
In this work, we have investigated the impact of trapped electron and muon neutrinos, muons and pions on the post-merger phase of a neutron star coalescence. Assuming that neutrinos are trapped and that the muons and pion decay are in equilibrium, we have constructed several equations of state systematically varying the microphysics content. We have then performed fully general-relativistic neutron star merger simulations to study the impact of these equations of state on the mass ejection and gravitational wave emission. Our results also present a detailed investigation into the regions of the remnant most sensitive to pion condensation.
@@
Although our results are part of a first number of steps towards a consistent inclusion of muon and muon neutrinos into neutron star merger simulations, they need to be improved in several ways. First, neutrinos should be modelled using fully consistent radiation transport schemes Foucart (2018); Radice et al. (2022); Izquierdo et al. (2024). This is particularly important in correctly tracking the muonic lepton fraction, which we have crudely approximated here. Furthermore, we treat $\pi$ as ideal Bose-Einstein gasses, while treating $\mu$ as ideal Fermi gasses. To be more consistent, modeling the transport of these charged particles would be ideal. We also neglect the density dependence of the pion mass. Similar to Vijayan et al. (2023), we use the vacuum mass to present optimistic estimates, to see if these effects are observable under ideal conditions.
For future work, we plan to also extend beyond the NSNS system and better quantify the influence of these particles in the core-collapse supernova context Bollig et al. (2017). Furthermore, how $\pi+\mu$ may couple with more exotic particles like axions remains an interesting route to pursue Diamond et al. (2024); Manzari et al. (2024).
@@ A.1 The Choice of Muonic Lepton Fraction $Y_{l,\mu}$
The choice of $Y_{l,\mu}$ requires some care. Informed by the results of Loffredo et al. (2023), the muonic lepton fraction at densities below $10^{14}$ g cm-3 has been shown to be trapped at small values. Following Appendix A in Vijayan et al. (2023)—and similar to the value used in Loffredo et al. (2023)—we fix $Y_{l,\mu}=0.01$, which has been shown to replicate muon lepton fractions in merger simulations.

### v2

@@ III.4 Gravitational Wave Emission
One of the main promising routes of obtaining additional information from the post-merger gravitational wave signal is via the dominant frequency peak of the gravitational wave strain Chatziioannou et al. (2017); Wijngaarden et al. (2022); Criswell et al. (2023); Breschi et al. (2022, 2023). In Fig. 15 we show the frequency spectrum of the post-merger gravitational wave strain, focusing mainly on the dominant frequency peak. We begin by comparing the impact of adding different microphysical correction to the equation of state. In line with previous works, we find that the adding of trapped electron neutrinos does not appreciable alter the gravitational wave emission Zappa et al. (2022); Most et al. (2024); Hammond et al. (2021), although current simulations are likely not in a convergent regime Espino et al. (2024); Most et al. (2024). We investigate the impact of adding $\pi$ to the EOS. We already found earlier that pions lead to a net softening (more compact neutron stars) of the equation of state, which generically should translate to higher frequencies. In line with this expectation, we find that for both EOS, adding pions causes a shift of around $100\,\rm Hz$ of the post-merger gravitational wave signal, although the impact on SFHo seems slightly more pronounced. Such difference is consistent with uncertainties of quasi-universal relations Vretinaris et al. (2020); Topolski et al. (2024). Finally, we also investigate the impact of $\mu$, and trapped $\nu_{\mu}$. Similar to the addition of trapped $\nu_{e}$, we find that the impact on the post-merger gravitational wave signal is small, though we caution that $\mu$ can in principle regulate the population of $\pi$, as seen in the amount of pion condensation (Fig. 5). While these frequencies are out of the current frequency sensitivity bands for the LIGO-Virgo-KAGRA detectors Barsotti et al. (2010), these calculations stress the importance of capturing proper microphysics for accurate predictions of future NSNS waveforms to optimize future high frequency GW detectors.
@@ IV Conclusions
In this work, we have investigated the impact of trapped electron and muon neutrinos, muons and pions on the post-merger phase of a neutron star coalescence. Assuming that neutrinos are trapped and that the muon and pion decay are in equilibrium, we have constructed several equations of state systematically varying the microphysics content. We have then performed fully general-relativistic neutron star merger simulations to study the impact of these equations of state on the mass ejection and gravitational wave emission. Our results also present a detailed investigation into the regions of the remnant most sensitive to pion condensation.
@@
Although our results are part of a first number of steps towards a consistent inclusion of muon and muon neutrinos into neutron star merger simulations, they need to be improved in several ways. First, neutrinos should be modelled using fully consistent radiation transport schemes Foucart (2018); Radice et al. (2022); Izquierdo et al. (2024); Ho-Yin Ng et al. (2024). This is particularly important in correctly tracking the muonic lepton fraction, which we have crudely approximated here. Furthermore, we treat $\pi$ as ideal Bose-Einstein gasses, while treating $\mu$ as ideal Fermi gasses. To be more consistent, modeling the transport of these charged particles would be ideal. We also neglect the density dependence of the pion mass. Similar to Vijayan et al. (2023), we use the vacuum mass to present optimistic estimates, to see if these effects are observable under ideal conditions.
For future work, we plan to also extend beyond the NSNS system and better quantify the influence of these particles in the core-collapse supernova context Bollig et al. (2017). Furthermore, how $\pi+\mu$ may couple with more exotic particles like axions remains an interesting route to pursue Diamond et al. (2024); Dev et al. (2024); Manzari et al. (2024).
@@ A.1 The Choice of Muonic Lepton Fraction $Y_{l,\mu}$
The choice of $Y_{l,\mu}$ requires some care. Informed by the results of Loffredo et al. (2023), the muonic lepton fraction at densities below $10^{14}$ g cm-3 has been shown to be trapped at small values. Following Appendix A in Vijayan et al. (2023)—and similar to the value used in Loffredo et al. (2023)—we fix $Y_{l,\mu}=0.01$, which has been shown to replicate muon lepton fractions in merger simulations. However, as the field evolves, more accurate approximations for $Y_{l,\mu}$ exist Gieg et al. (2024); Ho-Yin Ng et al. (2024).